1.State the amplitude and period of the trigonometric function f(x)= 3 sin x/2
2.State the amplitude and period of the trigonometric function f(x)= -1/2 sin (8x)
3. State the period and amplitude of the trigonometric function f(x)= 3/4 cos 12x

Answer by drk(1908)

1.State the amplitude and period of the trigonometric function f(x)= 3 sin x/2
amplitude = 3
period = 2pi/b = 2pi/(1/2) = 4pi
--
2.State the amplitude and period of the trigonometric function f(x)= -1/2 sin (8x)
amplitude = 1/2
period = 2pi/b = 2pi/8 = pi/4
--
3. State the period and amplitude of the trigonometric function f(x)= 3/4 cos 12x
amplitude = 3/4
period = 2pi/b = 2pi/12 = pi/6